A piano wire weighing 6.00 g and having a length of 90.0 cm emits a fundamental frequency corresponding to the “Middle C” (v = 261.63 Hz). Find the tension in the wire.

Grade:10

1 Answers

Navjyot Kalra
9 years ago
Sol. . I = 90 cm = 0.9 m m = (6/90) g/cm = (6/900) kg/mt f = 261.63 Hz f = 1/2√(T/m) ⇒ T = 1478.52 N = 1480 N.